OpenStudy (anonymous):

step by step what is this answer? 9-3 7/8=

Atsie (atsie):

9 - 3... 8, 7, 6 Subtract backwards three times. What do we do next?

OpenStudy (anonymous):

im sorry i dont understand but 9-3=6

Atsie (atsie):

It say's 9 - 3 so, if you count backwards three times on your finger's from 9, it will come to 6. Is that easier to understand?

OpenStudy (anonymous):

yes, sorry

Atsie (atsie):

Don't be sorry, we all have to learn sometime! :) Now out of curiosity, does 7/3 appear in a fraction form?

OpenStudy (anonymous):

no

Atsie (atsie):

Why is there no symbol in between the 9 - 3 and the 7/8 ? Is this how it looks in your book?

jhonyy9 (jhonyy9):

@Atsie i think this wan being 9 - 3 integer 7/8

OpenStudy (anonymous):

in my book it says 9-3 7/8

jhonyy9 (jhonyy9):

yes @Micheiou this make sens sure

Atsie (atsie):

Well, 9 - 3 is 6, so what are we supposed to use to get to the 7/8? @jhonyy9 Am I over thinking this? Help me out here.

jhonyy9 (jhonyy9):

ok. so first of all 3 7/8 = (3*8+7)/8 = (24+7)/8 = 31/8 and now you can continue it subtracting from 9

jhonyy9 (jhonyy9):

@Atsie do you can - like continue it for medal please ?

jhonyy9 (jhonyy9):

courage this is easy

jhonyy9 (jhonyy9):

@Atsie how you subtract from one integer a fraction please ?

Atsie (atsie):

Alright, subtracting 31/8 from 9 would be 5.12 in decimal form. @jhonyy9 medal received :)

jhonyy9 (jhonyy9):

ok. but please write your calculi here step by step to make the asker understanding wour way easy - thank you

jhonyy9 (jhonyy9):

your way - sorry my mistaque

Atsie (atsie):

31 / 8 is 3.8 now 3.8 - 9 would be 5.2

OpenStudy (anonymous):

I am not understanding this. The answer is going to be a fraction though

jhonyy9 (jhonyy9):

but there are 9 - 3 7/8 9 - 31/8 9 31 -- - --- common denominator is 8 so multiplie 9 by 8 1 8 9*8 -31 72 -31 41 ------- = --------- = ----- = ? 8 8 8 like step by step

Atsie (atsie):

ooh, I thought you were saying something else @jhonyy9 . Step by step that way.

jhonyy9 (jhonyy9):

@Micheiou hope you understand it now sure easy every steps how you can solving same exercises - yes ?

OpenStudy (anonymous):

Yes, thanks

jhonyy9 (jhonyy9):

@Atsie remember please we here on openstudy wan,like making asker understanding end learning step by step how can how need solving exercises in what they have called help

jhonyy9 (jhonyy9):

ok. @Micheiou was my pleasure good luck bye and anytime than i m here online